### v2

# GlycanML: A Multi-Task and Multi-Structure Benchmark for Glycan Machine Learning

Abstract
Glycans are basic biomolecules and perform essential functions within living organisms. The rapid increase of functional glycan data provides a good opportunity for machine learning solutions to glycan understanding. However, there still lacks a standard machine learning benchmark for glycan function prediction. In this work, we fill this blank by building a comprehensive benchmark for Glycan Machine Learning (GlycanML). The GlycanML benchmark consists of diverse types of tasks including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. Glycans can be represented by both sequences and graphs in GlycanML, which enables us to extensively evaluate sequence-based models and graph neural networks (GNNs) on benchmark tasks. Furthermore, by concurrently performing eight glycan taxonomy prediction tasks, we introduce the GlycanML-MTL testbed for multi-task learning (MTL) algorithms. Also, we evaluate how taxonomy prediction can boost other three function prediction tasks by MTL. Experimental results show the superiority of modeling glycans with multi-relational GNNs, and suitable MTL methods can further boost model performance. We provide all datasets and source codes at https://github.com/GlycanML/GlycanML and maintain a leaderboard at https://GlycanML.github.io/project.
1 Introduction
Glycans are fundamental biomolecules that play crucial roles in maintaining the normal physiological functions and health status of living organisms. They can regulate inflammatory responses [25], enable the recognition and communication between cells [67], preserve stable blood sugar levels [6], etc. Thanks to the advance of high-throughput sequencing techniques of glycans [65, 34], a large number of glycan data are accessible, e.g., the more than 240 thousand glycans stored in the GlyTouCan database [54]. This progress enables glycan function analysis by machine learning methods which are essentially data-driven.
There are some existing works that employ machine learning models to predict glycan taxonomy [11, 12], glycan immunogenicity [11, 12], glycosylation [42, 36] and protein-glycan interaction [40]. These works aim to solve one or several related glycan function prediction problems with either sequence models adapted from natural language processing (NLP) [11, 42] or graph neural networks (GNNs) [12, 40]. However, there still lacks a comprehensive benchmark studying the general effectiveness of various machine learning models on understanding diverse glycan functions, which hinders the progress of machine learning for glycan understanding. As a matter of fact, benchmark datasets greatly facilitate the machine learning research of other biomolecules like small molecules [27, 21] and proteins [46, 63].
Therefore, in this work, we take the initiative of building a Glycan Machine Learning (GlycanML) benchmark featured with diverse types of tasks and multiple glycan representation structures. The GlycanML benchmark consists of 11 benchmark tasks for understanding important glycan properties, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. For each task, we carefully split the benchmark dataset to evaluate the generalization ability of machine learning models in real-world scenarios. For example, in glycan taxonomy prediction, we leave out the glycans with unseen structural motifs during training for validation and test, which simulates the classification of newly discovered glycans in nature with novel molecular structures.
The GlycanML benchmark accommodates two glycan representation structures, i.e., glycan tokenized sequences and glycan planar graphs. For each structure, we adopt suitable machine learning models for representation learning, where sequence encoders such as CNN [24], LSTM [26] and Transformer [56] are employed for glycan sequence encoding, and both homogeneous GNNs [31, 57, 62] and heterogeneous GNNs [22, 48, 55] are used to encode glycan graphs. In addition, two state-of-the-art small molecule encoders [66, 45] are also studied for all-atom glycan modeling. We evaluate each model on all benchmark tasks to study its general effectiveness. The GlycanML benchmark also provides a testbed, namely GlycanML-MTL, for multi-task learning (MTL) algorithms, where an MTL method is asked to simultaneously solve eight glycan taxonomy prediction problems which are highly correlated. The performance on this testbed measures how well an MTL method can transfer the knowledge learned from different glycan taxonomies, e.g., transferring between species-level classification and genus-level classification. Taking a step further, we study the effect of MTL on boosting three function prediction tasks (i.e., immunogenicity, glycosylation and interaction prediction) by concurrently performing taxonomy prediction.
Benchmark results show that the RGCN model [48], a typical heterogeneous GNN, performs best on most benchmark tasks, and a simple two-layer CNN can surprisingly achieve competitive performance by using only condensed sequential information of glycan structures. The MTL methods with elaborate task-reweighting strategies can further enhance the performance of glycan taxonomy prediction, and learning the immunogenicity and glycosylation type of glycans jointly with their taxonomies also achieves benefits, showing the potential of MTL for glycan understanding. We hope the GlycanML benchmark will spark the interest of studying glycoscience with machine learning.
2 Related Work
Glycan machine learning. With the expanding size of experimental glycomics datasets, the integration of machine learning techniques into glycoinformatics shows considerable promise [8, 36]. Early approaches that adapt machine learning into glycomics research include the use of traditional ML algorithms (e.g., SVMs) to learn patterns from mass spectrometry data [33, 37], predict glycosylation sites [13, 35, 51, 44], and classify glycans [64]. Recently, thanks to the advancements in deep learning and new glycomics datasets, there has been a rise in studies applying deep learning to glycan and glycosylation modeling. DeepNGlyPred [42] seeks to identify N-glycosylated sequon from the N-GlyDE dataset [44]. SugarBase [10] is a comprehensive glycan database with metadata and analytical tools, and it supports the development of many glycan representation learning models, such as SweetOrigins [10], SweetTalk [9], SweetNet [12], glyBERT [18] and GNNGLY [1]. Deep learning-based methods like LectinOracle [40] and GlyNet [14] also made notable progress in predicting binding strengths between proteins and glycans.
However, there still lacks a comprehensive benchmark that incorporates diverse types of glycan understanding tasks and different glycan modeling methods like sequence-based and graph-based methods. Also, it is unknown how multi-task learning (MTL) influences the learning of glycan property prediction. In this work, we fill these blanks by introducing the GlycanML benchmark with multiple task types, multiple representation schemes of glycan structures, and an MTL testbed.
Biological machine learning benchmarks. To evaluate the performance of different machine learning methods in modeling biomolecules, it is necessary to establish large-scale standardized benchmarks. MoleculeNet [61] is a widely-used benchmark of small molecule modeling, evaluating the efficacy of both traditional machine learning and deep learning on predicting molecular properties. A recent work [70] proposes a benchmark for evaluating the performance of Large Language Models (LLMs) on molecular property prediction. In the field of protein modeling, the renowned CASP [32] competition is dedicated to establishing standards for protein structure prediction. Also, benchmark datasets are constructed for machine learning guided protein engineering [46, 19], protein design [20] and protein function annotation [63, 69]. Benchmark datasets are also established for other biomolecules like DNAs [28, 41] and RNAs [60]. In this work, we take the initiative of building a glycan machine learning benchmark for comprehensive glycan understanding.
3 Benchmark Tasks
The GlycanML benchmark consists of 11 benchmark tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction, as illustrated in Figure 1. We summarize the information of all tasks in Table 1.
Figure 1: Illustration of benchmark tasks. (a) Predicting the biological taxonomy of glycans at eight levels. (b) Judging whether a glycan is immunogenic or not in organisms. (c) Analyzing how a glycan glycosylates its target protein. (d) Given a protein and a glycan, predicting their binding affinity.
3.1 Glycan Taxonomy Prediction
Scientific significance. The taxonomy of glycans lays the foundation for glycomics research [4]. Biologists commonly classify glycans based on their origin under the hierarchical system of domain, kingdom, phylum, class, order, family, genus and species. Such a systematic classification helps us compare the similarities and differences between glycans, which further facilitates the study of glycan structures and functions. Also, the glycan taxonomy helps us understand the process of biological evolution. By comparing the structures of glycans in different organisms, we can infer their phylogenetic relationships and possible changes that may occur during evolution. Therefore, it could be very helpful to have an accurate glycan taxonomy predictor based on machine learning.
Task definition. In the GlycanML benchmark, we study glycan taxonomy prediction on domain, kingdom, phylum, class, order, family, genus and species levels, leading to eight individual tasks. These tasks are formulated as classification problems with 4, 11, 39, 101, 210, 415, 922 and 1,737 biological categories, respectively. Taking the class imbalance into consideration, we report Macro-F1 score for each task.
Benchmark dataset. We collect the glycans in the SugarBase database [10] that are fully annotated with domain, kingdom, phylum, class, order, family, genus and species labels, with 13,209 glycans in total. We then represent each glycan with the frequencies of popular motifs (i.e., those frequently occurring substructures in glycans), where the motif list proposed by Thomès et al. [53] is employed. Based on such representations, we cluster all glycans in the dataset by K-means ($K=10$), where 8 clusters are assigned to training, and the remaining two clusters are respectively utilized for validation and test. By using such dataset splits, this set of tasks evaluate how well a machine learning model can generalize across structurally distinct glycans.
Table 1: Benchmark task descriptions. We list each task along with its type, the average number of monosaccharides in each glycan for this task (in mean(std) format), dataset statistics, and evaluation metric. Abbr., Mono.: Monosaccharides.
Task Task type #Mono. per glycan #Sample #Train/Validation/Test Metric Taxonomy prediction of Domain Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Kingdom Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Phylum Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Class Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Order Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Family Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Genus Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Taxonomy prediction of Species Classification 6.39(3.51) 13,209 11,010/1,280/919 Macro-F1 Immunogenicity prediction Binary classification 7.30(3.78) 1,320 1,026/149/145 AUPRC Glycosylation type prediction Classification 9.04(3.96) 1,683 1,356/163/164 Macro-F1 Protein-Glycan interaction prediction Regression 6.56(4.54) 564,647 442,396/58,887/63,364 Spearman’s $\rho$
3.2 Glycan Immunogenicity Prediction
Scientific significance. Predicting the immunogenicity of glycans is of great significance for vaccine design and disease treatment. (1) Glycans are key components in many vaccines, especially in bacterial vaccines. By predicting the immunogenicity of glycans, researchers can design more effective vaccine formulations, thereby improving the protective effect of vaccines [29]. (2) In addition, certain glycans can inhibit tumor growth by activating the immune system [3], and therefore accurately predicting glycan immunogenicity can help optimize tumor treatment strategies.
Task definition. We formulate this task as a binary classification problem, i.e., predicting whether a glycan is immunogenic or not. We evaluate with the AUPRC metric to measure the trade-off between precision and recall of a model on immunogenic glycans.
Benchmark dataset. We select out all glycans in the SugarBase [10] whose immunogenicity is annotated based on evidences in literature, summing up to 1,320 glycans. As in glycan taxonomy prediction, we use the motif-based dataset splitting scheme to derive training, validation and test splits with an 8:1:1 ratio. In this way, we evaluate models’ generalization ability across structurally distinct glycans.
3.3 Glycosylation Type Prediction
Scientific significance. Glycans are a class of macromolecules with diverse biological activities, including immune system regulation, antitumor effects, antiviral effects, etc. By predicting the type of glycosylation, researchers can better understand the relationship between glycan structure and its functions. Understanding the structure-function relationship is crucial for designing and synthesizing glycan derivatives with specific biological activities [7].
Task definition. Given a glycan, we aim at predicting whether it forms N-glycosylation, O-glycosylation or maintains a free state, formulated as a three-way classification problem. The Macro-F1 score is used for evaluation.
Benchmark dataset. We traverse the GlyConnect database [2] and select out all glycans with glycosylation annotations, with 1,683 glycans in total. Upon these data, we again employ the motif-based dataset splitting scheme (introduced in Section 3.1) to construct training, validation and test splits with an 8:1:1 ratio. This task again assesses the generalization ability across the glycans with distinct structures.
3.4 Protein-Glycan Interaction Prediction
Scientific significance. The interactions between proteins and glycans play a crucial role in cellular signaling, affecting cell growth, differentiation, and apoptosis [58]. For example, glycans are one of the main components of the extracellular matrix (ECM), which interact with proteins such as collagen, laminin, and fibronectin to form the structural framework of ECM, providing physical support and passing biochemical signals to cells [17]. Understanding these interactions helps reveal how cells respond to external signals.
Task definition. Given a protein and a glycan, this task aims to regress their binding affinity, where the Z-score transformed relative fluorescence unit represents binding affinity. For this task, we adopt the Spearman’s correlation coefficient as the evaluation metric to measure how well a model ranks a set of protein-glycan pairs with different binding affinities.
Benchmark dataset. This benchmark dataset is built upon 564,647 protein-glycan interactions deposited in the LectinOracle database [40]. It is desired to have a model that can well generalize to new proteins against training ones, considering the continuous discovery of new proteins by sequencing techniques. Therefore, we split the dataset based on protein sequence similarity. Specifically, we first cluster all protein sequences using MMseqs2 [50] (minimum sequence identity within each cluster: 0.5), and then we derive training, validation and test proteins by splitting all clusters with an 8:1:1 ratio. Finally, samples of protein-glycan pairs are split according to the protein splits, deriving the benchmark dataset for protein-glycan interaction prediction. This splitting scheme is consistent with practical vaccine design applications where newly discovered lectins (proteins) could be used as the target for glycan binding.
4 Methods
Figure 2: Illustration of glycan representations. (a) The glycan tokenized sequence is derived by tokenizing the IUPAC-condensed sequence. (b) The glycan planar graph is constructed by transforming the IUPAC-condensed sequence to graph.
4.1 Representations
In the GlycanML benchmark, we adopt two glycan-specific representation structures, i.e., glycan tokenized sequence and glycan planar graph, as illustrated in Figure 2.
Glycan tokenized sequence. A glycan is commonly represented by an IUPAC-condensed sequence. For example, in the sequence “Glc(a1-4)Glc”, two glucoses are connected by an alpha-1,4-glycosidic bond, and this structure is the basic component of starch, a typical glycan. To process such sequences with machine learning models, a straightforward way is to tokenize the IUPAC-condensed sequence. Specifically, we regard each monosaccharide (e.g., Glc), each glycosidic bond (e.g., a1-4), and each bracket that indicates glycan branching (i.e., “[” and “]”) as a single token, which derives the glycan tokenized sequence, denoted as $x_{s}=\{s_{i}\}_{i=1}^{N}$. Various sequence encoders like Transformers [56] can then be applied to such tokenized sequences for glycan representation learning.
Glycan planar graph. Essentially, an IUPAC-condensed sequence describes the branching structure of a glycan, in which the part between brackets “[” and “]” denotes a side branch of the main branch, as illustrated in Figure 2. This structure is well represented by a planar graph $x_{g}=(\mathcal{V},\mathcal{E})$, in which nodes $\mathcal{V}$ denotes monosaccharides, and edges $\mathcal{E}$ denotes glycosidic bonds. In this way, graph neural networks (GNNs) are readily used for glycan modeling.
Table 2: Architectures of baseline models. Abbr., Params.: parameters; dim.: dimension; conv.: convolutional; attn.: attention; concat.: concatenate.
Model Input Layer Hidden Layers Output Layer #Params. Sequence encoders Shallow CNN [49] 128-dim. token embedding 2 $\times$ 1D conv. layers (hidden dim.: 128; max pooling over all tokens 191.7K kernel size: 5; stride: 1; padding: 2) ResNet [24] 512-dim. token embedding 8 $\times$ residual blocks (hidden dim.: 512; attentive weighted sum over all tokens 11.4M + 512-dim. positional embedding kernel size: 3; stride: 1; padding: 1) LSTM [26] 640-dim. token embedding 3 $\times$ bidirectional LSTM layers weighted sum over all tokens 26.7M (hidden dim.: 640) + linear (output dim.: 640) + Tanh Transformer [56] 512-dim. token embedding 4 $\times$ Transformer blocks (hidden dim.: 512; linear (output dim.: 512) + Tanh 21.4M + 512-dim. positional embedding #attn. heads: 8; activation: GELU) upon [CLS] token Homogeneous GNNs GCN [31] 128-dim. node embedding 3 $\times$ GCN layers concat. mean & max pooling 67.8K GAT [57] 128-dim. node embedding 3 $\times$ GAT layers (#attn. heads: 2) concat. mean & max pooling 69.4K GIN [62] 128-dim. node embedding 3 $\times$ GIN layers concat. mean & max pooling 117.4K Heterogeneous GNNs MPNN [22] 128-dim. node & edge embedding 3 $\times$ MPNN layers Set2Set pooling [59] (#steps: 3) 4.0M RGCN [48] 128-dim. node embedding 3 $\times$ RGCN layers concat. mean & max pooling 4.2M CompGCN [55] 128-dim. node embedding 3 $\times$ CompGCN layers concat. mean & max pooling 150.4K All-Atom Molecular Encoders Graphormer [66] 768-dim. node embedding 8 $\times$ Graphormer layers concat. mean & max pooling 28.5M GraphGPS [45] 512-dim. node & edge embedding 6 $\times$ GraphGPS layers concat. mean & max pooling 17.4M
4.2 Baselines
We include four types of models in our benchmark, i.e., sequence encoders for modeling glycan tokenized sequences, homogeneous and heterogeneous GNNs for modeling glycan planar graphs, and all-atom molecular encoders for modeling the all-atom molecular graphs of glycans, with 12 baseline models in total. Their detailed architectures are provided in Table 2.
Sequence encoders. We study the performance of four typical sequence encoders. Inspired by the success of shallow CNNs in modeling biological sequences like protein sequences [49, 63], we investigate (1) a 2-layer shallow CNN model along with (2) a deep residual network (ResNet) [24] with 8 hidden layers. These two CNN models mainly focus on capturing local information in glycan sequences. To investigate the importance of long context modeling for glycan understanding, we also include (3) a 3-layer bidirectional LSTM [26] and (4) a 4-layer Transformer encoder [56].
Homogeneous GNNs. Upon glycan planar graphs, standard GNNs designed for homogeneous graph modeling can be readily used to learn glycan representations. In our benchmark, three typical homogeneous GNNs, i.e., GCN [31], GAT [57] and GIN [62], serve as baselines, and they are all configured with 3 message passing layers.
Heterogeneous GNNs. As a matter of fact, modeling glycans as homogeneous graphs is suboptimal, in which the rich information within glycosidic bonds is fully ignored. To capture the complete information in glycan graphs, it is more proper to view them as heterogeneous graphs and employ heterogeneous GNNs for representation learning. Therefore, we adapt three popular heterogeneous GNNs, i.e., MPNN [22], RGCN [48] and CompGCN [55], to model glycan graphs, where each model is equipped with 3 message passing layers.
All-atom molecular encoders. Essentially, glycans are a kind of macromolecules with hundreds of atoms, and therefore we can directly employ all-atom molecular encoders designed for small molecules to model glycans. Among existing small molecule encoders, Graphormer [66] and GraphGPS [45] achieve prominent performance, and we include them in the proposed benchmark. We follow the default architecture of these two models in their original work, as shown in Table 2.
4.3 Model Pipelines
Depending on inputs, the benchmark tasks of GlycanML can be solved with two model pipelines.
Single-glycan prediction. This pipeline handles the tasks that predict the properties of individual glycans, including glycan taxonomy prediction, glycan immunogenicity prediction, and glycosylation type prediction. For each task, the glycan representation vector is first extracted by a glycan encoder and then passed to an MLP head for task-specific prediction.
Protein-glycan interaction prediction. Because of the additional input of protein, the protein-glycan interaction prediction task requires a different pipeline. Given a protein and a glycan, we first extract the protein representation vector with a protein encoder (e.g., the ESM-1b pre-trained protein language model [47] in this work) and extract the glycan representation vector with a glycan encoder, and these two vectors are then concatenated and sent to an MLP head for interaction prediction.
4.4 Multi-Task Learning
In GlycanML, the glycan taxonomy prediction tasks classify glycans under the hierarchical system from domain to species. These tasks are highly correlated and well-suited for multi-task learning (MTL) where related tasks are learned together for better generalization performance [68]. Therefore, we integrate eight glycan taxonomy prediction tasks in GlycanML as a testbed for MTL algorithms, named as the GlycanML-MTL benchmark.
On this benchmark, we study 6 representative MTL methods that focus on loss design and model optimization under the MTL setting. All these methods use the network architecture with hard parameter sharing [68], where all tasks share a common backbone encoder, and each task owns its individual prediction head. We introduce these methods below, with an abbreviation after each one.
  • Naive MTL (N-MTL): The most straightforward way to perform MTL is to sum up the losses of all tasks with equal weights and optimize the model with this loss summation. Denoting the losses of GlycanML-MTL tasks as $\mathcal{L}_{i}$ ($i=1,\cdots,8$), the naive MTL loss is defined as: $\mathcal{L}_{\mathrm{N-MTL}}=\sum_{i=1}^{8}\mathcal{L}_{i}$.
  • Gradient Normalization (GN) [15]: However, regarding all tasks equally is suboptimal, considering the varying difficulties of different tasks. Therefore, this method employs a weighted loss summation $\mathcal{L}_{\mathrm{GN}}=\sum_{i=1}^{8}w_{i}\mathcal{L}_{i}$, where the weights satisfy: $\sum_{i=1}^{8}w_{i}=8$. The main idea of gradient normalization is that different tasks should be trained at similar rates (i.e., similar speed of convergence). To achieve this goal, authors first deem the L2 norm of per-task gradient as the training rate of the task: $r_{i}=||\nabla_{\theta}w_{i}\mathcal{L}_{i}||_{2}$ ($\theta$ denotes model parameters), and all tasks are then pushed to have similar training rates by optimizing the loss $\mathcal{L}(w_{1},\cdots,w_{8})=\sum_{i=1}^{8}||r_{i}-\bar{r}||_{1}$ ($\bar{r}=(\sum_{i=1}^{8}r_{i})/8$). For each training step, this loss is first optimized w.r.t. loss weights $\{w_{i}\}_{i=1}^{8}$, and, using the updated loss weights, the MTL loss $\mathcal{L}_{\mathrm{GN}}$ optimizes whole model parameters.
  • Temperature Scaling (TS) [30]: For classification tasks, the sharpness of categorical distribution represents prediction uncertainty, further implying task difficulty. Inspired by this fact, the TS method seeks to weigh different tasks by scaling their classification logits. In this way, each task loss is defined as $\mathcal{L}^{\mathrm{TS}}_{i}=-\log(\mathrm{Softmax}(f_{\theta}(y|x)/\sigma_{i}^{2}))$ ($i=1,\cdots,8$), where $f_{\theta}(y|x)$ represents the classification logit of sample $x$ at class $y$, and $\sigma_{i}$ denotes the task-specific temperature parameter for scaling. The temperature parameters are learned along with the whole model.
  • Uncertainty Weighting (UW) [30]: Kendall et al. [30] shows that the temperature-scaled losses above can be approximated by a weighted summation of unscaled losses: $\mathcal{L}_{\mathrm{UW}}=\sum_{i=1}^{8}\mathcal{L}_{i}/\sigma_{i}^{2}+\log\sigma_{i}$, where the weighting parameters $\{\sigma_{i}\}_{i=1}^{8}$ are learnable. This method also weighs different tasks based on the uncertainty of task predictions.
  • Dynamic Weight Averaging (DWA) [39]: The loss scales along training can well indicate task convergence. Therefore, this method employs the ratio of consecutive losses to weigh different tasks: $w_{i}(t)=8\cdot\mathrm{Softmax}(\mathcal{L}_{i}(t)/\mathcal{L}_{i}(t-1))$, where $\mathcal{L}_{i}(t)$ denotes the loss of task $i$ at training step $t$. In this way, more weights are assigned to the tasks with slower convergence.
  • Dynamic Task Prioritization (DTP) [23]: This method maintains a key performance indicator (KPI) $\kappa_{i}(t)$ for each task along training (moving average of classification accuracy on our benchmark) and weighs different tasks in a focal loss [38] manner: $w_{i}(t)=-(1-\kappa_{i}(t))^{\gamma_{i}}\log\kappa_{i}(t)$, where $\gamma_{i}$ is the focusing hyperparameter for task $i$. Such a task reweighting scheme pays more attention to difficult tasks with low KPI.
5 Experiments
5.1 Experimental Setups
Model setups. For glycan taxonomy, immunogenicity and glycosylation type prediction tasks, upon the glycan embedding extracted by the glycan encoder, we use an MLP with 2 hidden layers and a ReLU nonlinearity in between to perform prediction. For protein-glycan interaction prediction, we use the ESM-1b pre-trained protein language model [47] to extract protein embedding, and, upon the concatenation of protein and glycan embeddings, the binding affinity is predicted by a 2-layer MLP with ReLU activation.
Training setups. We conduct every experiment on three seeds (0, 1 and 2) and report the mean and standard deviation of results. We train with an Adam optimizer (learning rate: $5\times 10^{-4}$, weight decay: $1\times 10^{-3}$) for 50 epochs on taxonomy, immunogenicity and glycosylation type prediction and for 10 epochs on interaction prediction. The batch size is set as 32 for interaction prediction and 256 for other tasks. For model training, we use cross entropy loss to train taxonomy and glycosylation type prediction tasks, use binary cross entropy loss to train immunogenicity prediction, and adopt mean squared error to train interaction prediction. For model selection, 10 times of validation are uniformly performed along the training process, and the checkpoint with the best validation performance is selected for test. For multi-task learning (MTL), the focusing parameter $\gamma$ of the dynamic task prioritization (DTP) method is set as 2.0, and the model selection of all MTL methods is based on the mean accuracy over all tasks on the validation set. We conduct all experiments on a local server with 100 CPU cores and 4 NVIDIA GeForce RTX 4090 GPUs (24GB). Our implementation is based on the PyTorch [43] deep learning library (BSD-style license) and TorchDrug [71] drug discovery platform (Apache-2.0 license).
Table 3: Benchmark results on single-task learning. We report mean (std) for each experiment. Three color scales of blue denote the first, second and third best performance. Abbr., Immuno: Immunogenicity; Glycos: Glycosylation.
Model Taxonomy Immuno (AUPRC) Glycos (Macro-F1) Interaction (Spearman’s $\rho$) Weighted Mean Rank Domain (Macro-F1) Kingdom (Macro-F1) Phylum (Macro-F1) Class (Macro-F1) Order (Macro-F1) Family (Macro-F1) Genus (Macro-F1) Species (Macro-F1) Sequence Encoders Shallow CNN 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.776(0.027) 0.898(0.009) 0.261(0.008) 4.97 ResNet 0.635(0.009) 0.505(0.025) 0.331(0.061) 0.301(0.010) 0.183(0.082) 0.165(0.019) 0.112(0.018) 0.073(0.007) 0.754(0.124) 0.919(0.004) 0.273(0.004) 4.63 LSTM 0.621(0.012) 0.566(0.076) 0.413(0.036) 0.272(0.029) 0.174(0.023) 0.145(0.012) 0.098(0.016) 0.078(0.008) 0.912(0.068) 0.862(0.016) 0.280(0.001) 4.71 Transformer 0.612(0.009) 0.546(0.079) 0.316(0.014) 0.235(0.022) 0.147(0.007) 0.114(0.039) 0.065(0.001) 0.047(0.008) 0.856(0.012) 0.729(0.069) 0.244(0.009) 8.06 Homogeneous GNNs GCN 0.635(0.001) 0.527(0.006) 0.325(0.024) 0.237(0.009) 0.147(0.005) 0.112(0.010) 0.095(0.009) 0.080(0.006) 0.688(0.023) 0.914(0.011) 0.233(0.009) 7.78 GAT 0.636(0.003) 0.523(0.007) 0.301(0.014) 0.265(0.012) 0.190(0.009) 0.130(0.005) 0.125(0.010) 0.103(0.009) 0.685(0.053) 0.934(0.038) 0.229(0.002) 7.22 GIN 0.632(0.004) 0.525(0.007) 0.322(0.046) 0.300(0.027) 0.179(0.002) 0.152(0.005) 0.116(0.022) 0.105(0.011) 0.716(0.051) 0.924(0.013) 0.249(0.004) 5.59 Heterogeneous GNNs MPNN 0.632(0.007) 0.638(0.050) 0.372(0.019) 0.326(0.015) 0.235(0.046) 0.161(0.004) 0.136(0.008) 0.104(0.009) 0.674(0.119) 0.910(0.006) 0.217(0.002) 8.09 RGCN 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.780(0.006) 0.948(0.004) 0.262(0.005) 2.19 CompGCN 0.629(0.004) 0.568(0.047) 0.410(0.013) 0.381(0.024) 0.226(0.011) 0.193(0.012) 0.166(0.009) 0.138(0.014) 0.692(0.006) 0.945(0.002) 0.257(0.004) 4.28 All-Atom Molecular Encoders Graphormer 0.640(0.006) 0.468(0.054) 0.249(0.041) 0.201(0.013) 0.142(0.019) 0.112(0.009) 0.077(0.006) 0.054(0.044) 0.637 (0.062) 0.856(0.009) 0.211(0.027) 11.09 GraphGPS 0.477(0.002) 0.511(0.040) 0.314(0.022) 0.261(0.051) 0.153(0.018) 0.134(0.008) 0.105(0.006) 0.065(0.017) 0.637 (0.075) 0.883(0.032) 0.247(0.016) 9.38
5.2 Benchmark Results on Single-Task Learning
In Table 3, we report the single-task performance of 12 representative glycan encoders on benchmark tasks. We measure the comprehensive performance of a model with its weighted mean rank over all tasks, where each taxonomy prediction task owns the weight of $1/8$ and each of the other three tasks owns the weight of 1, so as to balance between different types of tasks. Based on these results, we highlight the following findings:
  • Multi-relational GNNs show superiority in glycan modeling. Two typical multi-relational GNNs, i.e., RGCN and CompGCN, respectively rank first and second place in terms of weighted mean rank. Especially, RGCN achieves the best performance on 6 out of 11 benchmark tasks. Therefore, it is beneficial to model a glycan as a multi-relational graph, where different types of glycosidic bonds are deemed as different relations between monosaccharides.
  • A simple shallow CNN is surprisingly effective. It is observed that the 2-layer shallow CNN ranks among the top three on 5 out of 11 tasks, and it ranks fifth place in terms of weighted mean rank. We thereby demonstrate that such a shallow CNN model is sufficient to produce informative glycan representations and lead to competitive performance, which aligns with previous findings that shallow CNNs can well model biological sequences like protein sequences [49, 63].
  • It is important to utilize glycosidic bond information. We can observe clear performance gains of heterogeneous GNNs over homogeneous GNNs on glycan modeling, where in terms of weighted mean rank, three heterogeneous GNNs rank 1st, 2nd and 10th places, while three homogeneous GNNs rank 6th, 7th and 8th places. Compared to homogeneous GNNs that regard all glycosidic bonds as the same, heterogeneous GNNs fully utilize glycosidic bond information by individually treating each type of bonds, leading to obvious benefits.
  • Small molecule encoders can hardly handle glycan modeling. Though performing well on small molecule modeling, Graphormer and GraphGPS perform poorly on the GlycanML benchmark, where they rank last two places in terms of weighted mean rank. These results illustrate that the small molecule encoders originally designed to model tens of atoms cannot well model a macromolecule like a glycan with hundreds of atoms, calling for specifically designed models for all-atom glycan modeling.
Table 4: Benchmark results on multi-task learning. We report mean (std) for each experiment. Two color scales of blue denote the first and second best performance.
Method Domain Kingdom Phylum Class Order Family Genus Species Mean Macro-F1 Backbone encoder: Shallow CNN Single-Task 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.327(0.003) N-MTL 0.619(0.009) 0.549(0.040) 0.333(0.029) 0.344(0.015) 0.213(0.015) 0.167(0.014) 0.150(0.008) 0.121(0.010) 0.312(0.010) GN 0.646(0.043) 0.497(0.045) 0.344(0.064) 0.312(0.002) 0.233(0.042) 0.206(0.027) 0.175(0.038) 0.137(0.038) 0.319(0.012) TS 0.642(0.031) 0.609(0.059) 0.393(0.061) 0.355(0.013) 0.249(0.019) 0.187(0.009) 0.141(0.010) 0.113(0.009) 0.336(0.014) UW 0.632(0.005) 0.620(0.046) 0.409(0.025) 0.320(0.011) 0.230(0.012) 0.179(0.012) 0.152(0.001) 0.118(0.008) 0.332(0.005) DWA 0.619(0.012) 0.536(0.018) 0.354(0.013) 0.330(0.041) 0.218(0.011) 0.152(0.018) 0.138(0.015) 0.099(0.012) 0.305(0.013) DTP 0.633(0.004) 0.556(0.023) 0.341(0.041) 0.303(0.028) 0.213(0.034) 0.171(0.010) 0.145(0.011) 0.107(0.006) 0.309(0.017) Backbone encoder: RGCN Single-Task 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.360(0.009) N-MTL 0.641(0.001) 0.646(0.066) 0.402(0.042) 0.376(0.054) 0.231(0.025) 0.190(0.007) 0.154(0.018) 0.119(0.007) 0.345(0.020) GN 0.626(0.002) 0.612(0.050) 0.379(0.011) 0.364(0.017) 0.233(0.010) 0.169(0.006) 0.148(0.009) 0.119(0.006) 0.331(0.005) TS 0.640(0.003) 0.652(0.050) 0.461(0.013) 0.388(0.031) 0.267(0.024) 0.209(0.010) 0.174(0.002) 0.151(0.002) 0.368(0.016) UW 0.362(0.014) 0.568(0.045) 0.395(0.042) 0.386(0.045) 0.245(0.015) 0.199(0.009) 0.165(0.005) 0.133(0.012) 0.341(0.019) DWA 0.587(0.095) 0.616(0.112) 0.400(0.035) 0.379(0.024) 0.230(0.017) 0.177(0.023) 0.151(0.015) 0.110(0.005) 0.331(0.032) DTP 0.647(0.019) 0.637(0.083) 0.399(0.069) 0.370(0.023) 0.266(0.031) 0.228(0.039) 0.190(0.043) 0.158(0.035) 0.362(0.007)
5.3 Benchmark Results on Multi-Task Learning
In Table 4, we report the benchmark results of different MTL methods against single-task learning. We select shallow CNN and RGCN as the backbone encoder, and all MTL methods are evaluated on each of them. According to benchmark results, we have the findings below:
  • The temperature scaling (TS) approach performs best. On both shallow CNN and RGCN, the TS approach achieves the highest mean Macro-F1 score, and it outperforms single-task learning with a clear margin (i.e., 2.75% improvement in mean Macro-F1 score) when using shallow CNN as backbone encoder. Therefore, the TS approach can well balance the learning signals from different glycan taxonomy prediction tasks, leading to stable performance gains.
  • MTL methods are not always beneficial. On shallow CNN, only the TS and UW methods outperform single-task learning in terms of mean Macro-F1 score; on RGCN, only the TS and DTP methods outperform single-task learning in terms of mean Macro-F1 score. Actually, most MTL methods lead to performance decrease compared to single-task learning. These results suggest the high difficulty of balancing between different glycan taxonomy prediction tasks. More efforts are thus required to boost the MTL performance on the GlycanML-MTL testbed, which we leave as one of our major future works.
  • MTL is more helpful for difficult tasks on RGCN. From domain-level to species-level classification (i.e., from 4-way classification to 1,737-way classification), the task difficulty monotonically increases. Under RGCN, we can observe more benefits of MTL on difficult tasks. For example, the DTP method outperforms single-task learning on order-, family, genus- and species-level classification (i.e., tasks with more categories), while the DTP method is inferior on kingdom-, phylum- and class-level classification (i.e., tasks with fewer categories). Therefore, in MTL with RGCN, more attention is paid to difficult tasks, leading to better performance on these tasks.
Table 5: Benchmark results on multi-task learning for glycan function prediction. We report mean (std) for each experiment.
Method Immunogenicity (AUPRC) Glycosylation (Macro-F1) Interaction (Spearman’s $\rho$) Backbone encoder: Shallow CNN Single-Task 0.776(0.027) 0.898(0.009) 0.261(0.008) Multi-Task 0.792(0.028) 0.912(0.011) 0.257(0.002) Backbone encoder: RGCN Single-Task 0.780(0.006) 0.948(0.004) 0.262(0.005) Multi-Task 0.801(0.001) 0.963(0.009) 0.259(0.030)
5.4 The Effect of Multi-Task Learning on Glycan Function Prediction
In this part, we study the effect of MTL on three glycan function prediction tasks, i.e., immunogenicity, glycosylation and interaction prediction. Considering that evolutionarily similar glycans tend to have similar functions, we study how glycan function prediction can be benefited by jointly learning the evolutionary taxonomy of glycans. Specifically, we regard each of three function prediction tasks as the center task and eight taxonomy prediction tasks as auxiliary tasks, and the center task is trained along with eight auxiliary tasks to perform MTL (the losses of center and auxiliary tasks are summed up for model training).
In Table 5, we compare this multi-task method with the single-task baseline. Under both Shallow CNN and RGCN encoders, MTL performs better on immunogenicity and glycosylation type prediction, while single-task learning performs better on interaction prediction. These results show the potential of MTL on boosting glycan function prediction, and also inspire future research efforts to further improve its effectiveness.
6 Conclusions and Future Work
In this work, we build a comprehensive benchmark GlycanML for glycan machine learning. It consists of diverse types of glycan understanding tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. In GlycanML, we support two representation methods of glycans, i.e., glycan tokenized sequences and glycan planar graphs. Additionally, on eight highly correlated glycan taxonomy prediction tasks, we set up a testbed GlycanML-MTL to compare different multi-task learning (MTL) algorithms. Also, we study how taxonomy prediction can boost other three function prediction tasks by MTL. According to the benchmark results, multi-relational graph neural networks (GNNs) show great promise for glycan modeling, and well-designed MTL methods can further boost model performance.
Limitations and future work. The current GlycanML benchmark is limited to modeling glycan sequences and 2D glycan graphs, without benchmark datasets of 3D glycan structures. Therefore, in the future, we will go beyond sequence- and 2D-graph-based datasets and methods to 3D-structure-based ones [5, 16]. Also, we will work along with the community to further promote the efforts of MTL for better glycan understanding.
7 Broader Societal Impacts
This work aims at building a comprehensive benchmark for glycan machine learning, incorporating diverse types of glycan understanding tasks, multiple representation methods of glycan structures, and a testbed for multi-task learning algorithms. By evaluating on the proposed benchmark, we can judge the general effectiveness of a machine learning model on predicting different glycan functions. Therefore, such a benchmark can promote the application of machine learning methods to various real-world glycan-related tasks, such as vaccine design [29] and cancer research [52].
However, we should not ignore the potential harmful aspects brought by glycan machine learning models developed on our benchmark, e.g., designing vaccines with severe adverse reactions. To mitigate such risks, our future works will encourage the responsible usage of the GlycanML benchmark for real-world problems.